OpenStudy (anonymous):

log2(3/8)-log2(3), probably easy but logs are not my thing.. help?

OpenStudy (anonymous):

= log2(3)-log2(8)-log2(3) = -log2(8)

OpenStudy (anonymous):

is it log(base2)(3/8)-log(base2)(3)?

OpenStudy (anonymous):

yeah, definitively

OpenStudy (anonymous):

= log2(3)-log2(8)-log2(3) = -log2(2^3) =-3 Sorry I missed this last step.

OpenStudy (anonymous):

thank you so much! i am starting to like 'em logs now

OpenStudy (anonymous):

Thats good. Dont forget to click the "Good Answer" buttom

OpenStudy (anonymous):

any time :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

log(base2)(3/8)-log(base2)(3) =log(base2)(3/8*3) =log(base2)(1/8) =log(base2)(1)-log(base2)(8) =0-log(base2)(8) =-log(base2)(8) ans...

OpenStudy (anonymous):

yeah, but -log2(8) = -3 this is a longer way